OpenStudy (anonymous):

Derivative of [sin [(x+1)/(x-1)]]^3? show steps please

OpenStudy (anonymous):

you need to use the chain rule several times like let u = (x + 1)/(x -1) so you have sin^3(u) then let v = sin u and take the derivative of v^3 back substitute everything in its long, but doable

OpenStudy (anonymous):

I have 3[sin (x+1)/(x-1)]^2 times cos(x+10/(x-1)

OpenStudy (anonymous):

not sure. try using wolframalpha to verify if that is correct. what you get should be equivalent to \[-\frac{ 3 \sin \left( \frac{ x+1 }{ x - 1 } \sin \left( \frac{ 2(x+1) }{ x - 1 } \right) \right) }{ (x - 1)^{2} }\]